# Benchmarking MedMNIST dataset on real quantum hardware

### v1

@@ Abstract
Quantum machine learning (QML) has emerged as a promising domain to leverage the computational capabilities of quantum systems to solve complex classification tasks. In this work, we present first comprehensive QML study by benchmarking the MedMNIST-a diverse collection of medical imaging datasets on a 127-qubit real IBM quantum hardware, to evaluate the feasibility and performance of quantum models (without any classical neural networks) in practical applications. This study explore recent advancements in quantum computing such as device-aware quantum circuits, error suppression and mitigation for medical image classification. Our methodology comprised of three stages: preprocessing, generation of noise-resilient and hardware-efficient quantum circuits, optimizing/training of quantum circuits on classical hardware, and inference on real IBM quantum hardware. Firstly, we process all input images in the preprocessing stage to reduce the spatial dimension due to the quantum hardware limitations. We generate hardware-efficient quantum circuits using backend properties expressible to learn complex patterns for medical image classification. After classical optimization of QML models, we perform the inference on real quantum hardware. We also incorporates advanced error suppression and mitigation techniques in our QML workflow including dynamical decoupling (DD), gate twirling, and matrix-free measurement mitigation (M3) to mitigate the effects of noise and improve classification performance. The experimental results showcase the potential of quantum computing for medical imaging and establishes a benchmark for future advancements in QML applied to healthcare.
@@ 1 Introduction
Biomedical image analysis plays a critical role in advancing human healthcare. Recently, machine learning (ML) has demonstrated remarkable potential in this domain, excelling in tasks such as medical image classification [1, 2, 3], segmentation [4, 5], and reconstruction [6, 7]. Simultaneously, QML an emerging paradigm of ML implemented on quantum computers, is anticipated to offer quantum advantages over classical ML. These advantages include enhanced speed-ups, improved model expressivity, efficient exploration of high-dimensional feature spaces, and the ability to address certain classically intractable problems [8, 9, 10, 11, 12]. QML capitalizes on the unique properties of quantum mechanics, such as superposition, entanglement, and interference, to unravel complex relationships between inputs and labels. In the current noisy intermediate-scale quantum (NISQ) era, QML faces notable challenges. Limited access to large-scale quantum systems constrains the ability to encode high-dimensional data into quantum states. Hardware noise and decoherence further impact accuracy and restrict circuit depth, thereby limiting the expressivity of QML models. Rapid advancements in quantum hardware, including improved quantum error correction techniques [13], are expected to overcome these limitations and enable the practical deployment of QML for real-world applications in the near future.
@@
Despite these obstacles, QML research has witnessed significant growth [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Generally, QML approaches are based on parameterized quantum circuits (PQCs) [29] consisting of fixed gates (such as SWAP and CNOT) and single-qubit (such as Pauli rotation) gates with trainable parameters. These PQCs comprised of three stages: data encoding, a variational ansatz circuit with trainable parameters, and measurement to classical bits. The data encoding stage encode the classical input data to quantum states. The parameters of the ansatz circuit are optimized using classical gradient descent algorithms, typically relying on the parameter-shift rule [30], enabling the circuit to solve machine learning tasks.
@@
Over the last few years, researchers have explored QML algorithms for medical imaging in various studies including [26, 31, 32, 33, 34, 35, 36, 37, 38, 39] to harness potential quantum advantages over classical machine learning approaches. These studies address a variety of medical imaging modalities, including MRI, CT, X-ray, OCT, and ultrasound, targeting tasks such as clinical data classification and detection. The QML algorithms investigated in these works include quantum support vector machines (SVM) [34], hybrid quantum-classical neural networks [26, 35, 36, 38, 39], quantum transfer learning [33, 37], and variational quantum circuits [31, 32]. To the best of our knowledge, few studies such as [26, 33, 34, 37, 39] have used real quantum hardware for evaluation purposes but at a limited scale. Cherrat et al. proposed a quantum vision transformer (QViT) in [26] based on shallow PQCs capable of performing complex image classification tasks. This QViT introduces trainable quantum orthogonal layers adaptable to varying levels of quantum hardware connectivity. The authors have evaluated their QViT on open source MedMNIST (a collection of 12 biomedical imaging datasets) [1] using a simulator and real hardware results are provided for only one dataset (RetinaMNIST) out of 12. In [33], Acar and Yilmaz designed a hybrid classical-quantum transfer learning based QML model for COVID-19 detection using CT images. The authors claimed to achieve a classification accuracy of 94-100% on quantum devices with limited qubit capacity compared to 90% on classical systems using a small dataset. This highlights the potential of quantum computers for efficient and accurate medical image classification, particularly with limited data. Moradi et al. [34] designed two QML methods i.e., quantum distance classifier and quantum-kernel SVM for clinical data classification on quantum hardware. The findings observed in this study also highlight the promise of QML in clinical data analysis, especially for tasks constrained by dataset size and feature complexity. Azevedo et al. [37] employed hybrid quantum-classical neural network for mammogram classification into malignant and benign categories. The authors compared the results obtained from quantum device and simulator, highlighting the potential of QML in medical diagnostics while emphasizing the need for further validation. Similarly, the study conducted in [39] by Landman et al. introduces two quantum neural networks: quantum-assisted neural networks, which leverage quantum computer to estimate inner products during the training and inference of classical neural networks, and quantum orthogonal neural networks, built on quantum pyramidal circuits for orthogonal matrix multiplication. This study also considered only two datasets (PneumoniaMNIST and RetinaMnist) out of 12 datasets from MedMNIST for evaluation purposes on real hardware. Moreover, all of these studies showing evaluation on quantum hardwares are small-scale devices with 5, 7, 15, 16, and 27 qubits. It is also worth noting that most of these QML studies are based on hybrid quantum-classical neural networks which combines quantum circuits with classical neural networks, leaving the potential of purely quantum approaches underexplored. Also, the individual contribution of both quantum and classical parts is unclear. While some studies claim potential quantum advantages, these are often observed only in restricted settings (e.g., small datasets or low-qubit systems) with limited evidence of their generalization across different image classification domains or modalities. Demonstrating clear, reproducible quantum advantages over classical machine learning models remains a significant challenge.
@@
  • We present the first comprehensive QML study by benchmarking the MedMNIST dataset on real IBM Cleveland quantum hardware.
@@
In the remaining sections of this research study, we firstly discussed our QML methodology in detail in Section 2. Then, we provide the experiment results with discussion in Section 3. Lastly, concluding remarks are provided in Section 4.
@@ 2 Methodology
In this section, we discuss our QML methodology for medical image classification, as illustrated in Fig. 1. Our approach is structured into four key stages: data preprocessing, generation of device-aware quantum circuits, training of the quantum circuits on classical hardware using training dataset, and inference of the trained quantum model on real IBM quantum hardware using a testing dataset. It is important to emphasize that, due to the limitations of current NISQ quantum devices, training quantum models with large datasets directly on hardware remains infeasible. To address this challenge, our QML workflow adopts a hybrid strategy, where the training phase is executed on classical hardware, leveraging their computational power, while the inference stage is performed on real quantum hardware to validate the model’s performance. The following subsections provide a detailed explanation of each stage of our QML methodology.
@@ 2.1 Data preprocessing
The preprocessing stage focuses on reducing the feature dimensionality of input images, a necessary step in our QML workflow due to the constraints of current quantum hardware. Specifically, we employ angle encoding to map classical data to quantum states. The angle encoding method is limited by the hardware’s inability to handle high-dimensional feature spaces in the current NISQ era. Therefore, it is very important to reduce the spatial dimension of the images effectively. We used average pooling to downscale the input image dimension from 224$\times$224 to 7$\times$7 and 8$\times$8 for the considered MedMNIST datasets, thereby providing 49 and 64 features, respectively. This indicates that performing angle encoding for 49 and 64 features requires the incorporation of 49 and 64 Pauli rotation gates, respectively, within the quantum circuit. The quantum circuit representational capacity should be adjusted accordingly to effectively capture the patterns in the input features. Post-feature reduction, the pixel values are normalized to the range $[0,\pi]$, aligning with the requirements of angle encoding for efficient quantum state preparation. It is important to note that this preprocessing stage introduces information loss compared to classical machine learning models, which can process entire images as input without dimensionality reduction. In QML workflow, this limitation arises due to the constraints of current quantum hardware, which restrict the number of features that can be encoded into quantum states. Consequently, while classical models can leverage the full spatial resolution of input images, quantum models operating with a reduced feature set, potentially impacting the overall representational capacity. This trade-off underscores the need for effective feature reduction techniques that preserve essential information while ensuring compatibility with the limitations of quantum devices.
@@ 2.2 Efficient and noise-resilient quantum circuits for QML
The second crucial step of our QML framework involves designing a noise-robust and efficient quantum circuit that is expressively powerful enough to learn meaningful patterns from the input data for medical imaging classification task. In the current NISQ era, developing device-aware quantum circuits tailored to the target backend properties, such as qubit connectivity and gate fidelities, has emerged as a key strategy to mitigate the impact of hardware noise. However, designing such circuits remains a significant challenge due to the lack of well-defined guidelines for quantum circuit architecture in QML workflows. Additionally, the optimal design of a quantum circuit is highly dataset-dependent, as varying datasets require different levels of expressivity to effectively capture their underlying patterns.
@@
To address these challenges, we incorporated the Élivágar framework [40] into our QML workflow to generate hardware-aware quantum circuits tailored to our medical image classification task. This framework allowed us to create noise-resilient and computationally efficient quantum circuits that align with the physical constraints of the target quantum hardware while maximizing model performance. The quantum circuits in our workflow comprise three key components: (1) data encoding gates, (2) a variational or ansatz circuit, and (3) measurement operations that map quantum states to classical bits. The data encoding layer leverages parameterized rotation gates to embed classical input data into quantum states. The ansatz, or variational circuit, contains trainable Pauli rotation gates interspersed with CNOT gates to ensure qubit entanglement. The parameters of these gates ($\theta$) are optimized during training to minimize the loss function. Input data from the preprocessing stage is encoded using a number of rotation gates equal to the feature count of each input sample. An essential step in the Élivágar framework is computing the Clifford Noise Resilience (CNR) to assess the robustness of candidate circuits against hardware-induced errors. While this process is typically conducted on quantum device, it is time-intensive due to the need to run thousands of circuits, often facing significant queuing delays. To streamline the process, we performed CNR calculations using a simulator, enabling efficient evaluation of circuit noise resilience while saving considerable time.
@@ 2.3 Training quantum model on classical hardware
The training phase of our QML framework for medical image classification is executed on classical hardware as illustrated in Fig. 1. This approach circumvents the current limitations of quantum hardware in the NISQ era, such as limited qubits, high noise levels, and decoherence effects by performing the computationally expensive training process on classical GPUs even for large-scale datasets. Our QML workflow integrates classical optimization technique with noise-resilient, hardware-efficient quantum circuits tailored for medical imaging tasks. The training pipeline includes iterative optimization using gradient-based method and loss evaluation. The variational quantum circuit parameters $(\theta)$ are iteratively updated within a classical optimization loop to minimize a task-specific loss function. This iterative process continues until convergence is achieved or the maximum number of training epochs is reached, ensuring the parameters are optimized for the target classification task.
We conducted the training on noiseless simulators using the TorchQuantum library [41], which provides an ideal environment for parameter optimization without interference from hardware-induced noise. This approach significantly reduces computational overhead compared to direct hardware training and ensures consistent gradient computation. During each training iteration, the preprocessed input data is mapped to quantum states via encoding gates in the quantum circuit, which is then executed on the simulator. Expectation values obtained from circuit measurements are used to compute the loss function. Gradients of the loss function are calculated using the parameter-shift rule [30]. Afterwards, these gradients guide precise updates to the trainable parameters $(\theta)$. In our implementation, all the quantum circuits were trained for 200 epochs using the Adam optimizer with a learning rate of 0.01 and a batch size of 128, ensuring convergence to optimal parameter configurations. The MedMNIST contains datasets having different number of classes. During training, we measure one qubit and two qubits for two class and four class datasets, respectively and mean square loss function is used to calculate the loss. For the remaining multi-class datasets, we measure the number of qubits equals to the number of classes in the dataset and cross entropy loss function is used to calculate the loss. To ensure compatibility with MedMNIST datasets, we trained circuits with varying numbers of trainable parameters, such as 60, 80, 100, and 120 for each dataset of MedMNIST to account for differences in dataset complexity. The best-performing models are saved along with their parameters, losses, and accuracies for later inference on real quantum hardware. We also performed inference of these trained quantum models on simulator for all the considered MedMNIST datasets. All the quantum models are trained on a A100 Nvidia GPU of 43 GB RAM.

### v2

@@ Abstract
Quantum machine learning (QML) has emerged as a promising domain to leverage the computational capabilities of quantum systems to solve complex classification tasks. In this work, we present the first comprehensive QML study by benchmarking the MedMNIST-a diverse collection of medical imaging datasets on a 127-qubit real IBM quantum hardware, to evaluate the feasibility and performance of quantum models (without any classical neural networks) in practical applications. This study explores recent advancements in quantum computing such as device-aware quantum circuits, error suppression, and mitigation for medical image classification. Our methodology is comprised of three stages: preprocessing, generation of noise-resilient and hardware-efficient quantum circuits, optimizing/training of quantum circuits on classical hardware, and inference on real IBM quantum hardware. Firstly, we process all input images in the preprocessing stage to reduce the spatial dimension due to quantum hardware limitations. We generate hardware-efficient quantum circuits using backend properties expressible to learn complex patterns for medical image classification. After classical optimization of QML models, we perform inference on real quantum hardware. We also incorporate advanced error suppression and mitigation techniques in our QML workflow, including dynamical decoupling (DD), gate twirling, and matrix-free measurement mitigation (M3) to mitigate the effects of noise and improve classification performance. The experimental results showcase the potential of quantum computing for medical imaging and establish a benchmark for future advancements in QML applied to healthcare.
@@ 1 Introduction
Biomedical image analysis plays a critical role in advancing human healthcare. Recently, machine learning (ML) has demonstrated remarkable potential in this domain, excelling in tasks such as medical image classification [1, 2, 3], segmentation [4, 5], and reconstruction [6, 7]. Moreover, QML, an emerging paradigm of ML implemented on quantum computers, is anticipated to offer quantum advantages over classical ML. These advantages include enhanced speed-ups, improved model expressivity, efficient exploration of high-dimensional feature spaces, and the ability to address certain classically intractable problems [8, 9, 10, 11, 12]. QML capitalizes on the unique properties of quantum mechanics, such as superposition, entanglement, and interference, to unravel complex relationships between inputs and labels. In the current noisy intermediate-scale quantum (NISQ) era, QML faces notable challenges. Limited access to large-scale quantum systems constrains the ability to encode high-dimensional data into quantum states. Hardware noise and decoherence further impact accuracy and restrict circuit depth, thereby limiting the expressivity of QML models. Rapid advancements in quantum hardware, including improved quantum error correction techniques [13], are expected to overcome these limitations and enable the practical deployment of QML for real-world applications in the near future.
@@
Despite these obstacles, QML research has witnessed significant growth [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Generally, QML approaches are based on parameterized quantum circuits (PQCs) [29] consisting of fixed gates (such as SWAP and CNOT) and single-qubit (such as Pauli rotation) gates with trainable parameters. These PQCs are comprised of three stages: data encoding, a variational ansatz circuit with trainable parameters, and measurement to classical bits. The data encoding stage encodes the classical input data to quantum states. The parameters of the ansatz circuit are optimized using classical gradient descent algorithms, typically relying on the parameter-shift rule [30], enabling the circuit to solve machine learning tasks.
@@
Over the last few years, researchers have explored QML algorithms for medical imaging in various studies including [26, 31, 32, 33, 34, 35, 36, 37, 38, 39] to harness potential quantum advantages over classical machine learning approaches. These studies address a variety of medical imaging modalities, including MRI, CT, X-ray, OCT, and ultrasound, targeting tasks such as clinical data classification and detection. The QML algorithms investigated in these works include quantum support vector machines (SVM) [34], hybrid quantum-classical neural networks [26, 35, 36, 38, 39], quantum transfer learning [33, 37], and variational quantum circuits [31, 32]. To the best of our knowledge, few studies such as [26, 33, 34, 37, 39] have used real quantum hardware for evaluation purposes but on a limited scale. Cherrat et al. proposed a quantum vision transformer (QViT) in [26] based on shallow PQCs capable of performing complex image classification tasks. QViT introduces trainable quantum orthogonal layers adaptable to varying levels of quantum hardware connectivity. The authors have evaluated QViT on open-source MedMNIST (a collection of 12 biomedical imaging datasets) [1] using a simulator and real hardware results are provided for only one dataset (RetinaMNIST) out of 12 total. In [33], Acar and Yilmaz designed a hybrid classical-quantum transfer learning based QML model for COVID-19 detection using CT images. The authors claimed to achieve a classification accuracy of 94-100% on quantum devices with limited qubit capacity compared to 90% on classical systems using a small dataset. This highlights the potential of quantum computers for efficient and accurate medical image classification, particularly with limited data. Moradi et al. [34] designed two QML methods i.e., quantum distance classifier and quantum-kernel SVM for clinical data classification on quantum hardware. The findings observed in this study also highlight the promise of QML in clinical data analysis, especially for tasks constrained by dataset size and feature complexity. Azevedo et al. [37] employed a hybrid quantum-classical neural network for mammogram classification into malignant and benign categories. The authors compared the results obtained from the quantum device and simulator, highlighting the potential of QML in medical diagnostics while emphasizing the need for further validation. Similarly, the study conducted in [39] by Landman et al. introduces two quantum neural networks: quantum-assisted neural networks, which leverage the quantum computer to estimate inner products during the training and inference of classical neural networks, and quantum orthogonal neural networks, built on quantum pyramidal circuits for orthogonal matrix multiplication. This study also considered only two datasets (PneumoniaMNIST and RetinaMNIST) out of 12 datasets from MedMNIST for evaluation purposes on real hardware. It is also worth noting that most of these QML studies are based on hybrid quantum-classical neural networks that combine quantum circuits with classical neural networks, leaving the potential of purely quantum approaches underexplored. Also, the individual contribution of both quantum and classical parts is unclear. While some studies claim potential quantum advantages, these are often observed only in restricted settings (e.g., small datasets) with limited evidence of their generalization across different image classification domains or modalities. Demonstrating clear, reproducible quantum advantages over classical machine learning models remains a significant challenge.
@@
  • We present the first comprehensive QML study by benchmarking the MedMNIST dataset on IBM Cleveland.
@@
In the remaining sections of this research study, we first discussed our QML methodology in detail in Section 2. Then, we provide the experimental results with detailed discussion in Section 3. Lastly, concluding remarks are provided in Section 4.
@@ 2 Methodology
In this section, we discuss our QML methodology for medical image classification, as illustrated in Fig. 1. Our approach is structured into four key stages: data preprocessing, generation of device-aware quantum circuits, training of the quantum circuits on classical hardware using a training dataset, and inference of the trained quantum model on real IBM quantum hardware using a testing dataset. It is important to emphasize that, due to the limitations of current NISQ quantum devices, training quantum models with large datasets directly on hardware remains infeasible. To address this challenge, our QML workflow adopts a hybrid strategy, where the training phase is executed on classical hardware, leveraging their computational power, while the inference stage is performed on real quantum hardware to validate the model’s performance. The following subsections provide a detailed explanation of each stage of our QML methodology.
@@ 2.1 Data preprocessing
The preprocessing stage focuses on reducing the feature dimensionality of input images, a necessary step in our QML workflow due to the constraints of current quantum hardware. Specifically, we employ angle encoding to map classical data to quantum states. The angle encoding method is limited by the hardware’s inability to handle high-dimensional feature spaces in the current NISQ era. Therefore, it is very important to reduce the spatial dimension of the images effectively. We used average pooling to downscale the input image dimension from 224$\times$224 to 7$\times$7 and 8$\times$8 for the considered MedMNIST datasets, thereby providing 49 and 64 features, respectively. This indicates that performing angle encoding for 49 and 64 features requires the incorporation of 49 and 64 Pauli rotation gates, respectively, within the quantum circuit. The quantum circuit representational capacity should be adjusted accordingly to effectively capture the patterns in the input features. Post-feature reduction, the pixel values are normalized to the range $[0,\pi]$, aligning with the requirements of angle encoding for efficient quantum state preparation. It is important to note that this preprocessing stage introduces information loss compared to classical machine learning models, which can process entire images as input without dimensionality reduction. In a QML workflow, this limitation arises due to the constraints of current quantum hardware, which restrict the number of features that can be encoded into quantum states. Consequently, while classical models can leverage the full spatial resolution of input images, quantum models operate with a reduced feature set, potentially impacting the overall representational capacity. This trade-off underscores the need for effective feature-reduction techniques that preserve essential information while ensuring compatibility with the limitations of quantum devices.
@@ 2.2 Efficient and noise-resilient quantum circuits for QML
The second crucial step of our QML framework involves designing a noise-robust and efficient quantum circuit that is expressively powerful enough to learn meaningful patterns from the input data for the medical imaging classification tasks. In the current NISQ era, developing device-aware quantum circuits tailored to the target backend properties, such as qubit connectivity and gate fidelities, has emerged as a key strategy to mitigate the impact of hardware noise. However, designing such circuits remains a significant challenge due to the lack of well-defined guidelines for quantum circuit architecture in QML workflows. Additionally, the optimal design of a quantum circuit is highly dataset-dependent, as varying datasets require different levels of expressivity to effectively capture their underlying patterns.
@@
To address these challenges, we incorporated the Élivágar framework [40] into our QML workflow to generate hardware-aware quantum circuits tailored to our medical image classification task. This framework allowed us to create noise-resilient and computationally efficient quantum circuits that align with the physical constraints of the target quantum hardware while maximizing model performance. The first step of the Élivágar framework is to generate candidate quantum circuits using information from the target quantum device. We generated 250 quantum circuits in all the MedMNIST dataset experiments. These quantum circuits include three key components: (1) data encoding gates, (2) a variational or ansatz circuit, and (3) measurement operations that map quantum states to classical bits. The data encoding layer leverages parameterized rotation gates to embed classical input data into quantum states. The ansatz, or variational circuit, contains trainable Pauli rotation gates interspersed with CNOT gates to ensure qubit entanglement. The parameters of these gates ($\theta$) are optimized during training to minimize the loss function. Input data from the preprocessing stage is encoded using some rotation gates equal to the feature count of each input sample. Then, Clifford Noise Resilience (CNR) [40] is computed to assess the robustness of these candidate quantum circuits against hardware noise. This process involves the generation and execution of multiple Clifford replicas of the circuit ($C$) denoted by $C_{R}(i)$, followed by a fidelity assessment between the noisy ($P_{noisy}$) and noiseless ($P_{noiseless}$) outputs. To achieve this, the Total Variation Distance (TVD) is first evaluated between noiseless and noise outputs as defined in the Eq. (1) and the circuit fidelity $(F)$ is computed by $F=1-TVD$.
$TVD=\frac{1}{2}\sum_{i\in[2^{n_{q}}]}\left|P_{noiseless}(|i\rangle)-P_{noisy}(|i\rangle)\right|$ (1)
where $n_{q}$ denotes the number of qubits in the circuit. The CNR of a circuit is then computed as the average fidelity of the $M$ Clifford replicas defined as:
$CNR(C)=\frac{1}{M}\sum_{i=1}^{M}F(P_{noiseless}^{C_{R}(i)},P_{noisy}^{C_{R}(i)})$ (2)
The low-fidelity candidate circuits with CNR values lower than the threshold value (0.7) are excluded from further evaluation. While this process is typically conducted on a quantum device, it is time-intensive due to the need to run thousands of circuits, often facing significant queuing delays. To streamline the process, we performed CNR calculations using IBM Cleveland’s hardware noise model on a simulator (with the number of shots = 10000 and $M$ = 32), enabling efficient evaluation of circuit noise resilience while saving considerable time. Afterward, the performance of selected quantum circuits is assessed using Representation Capacity (RepCap) [40] defined as:
$RepCap(C)=1-\frac{\|R_{C}-R_{\text{ref}}\|_{2}^{2}}{2n_{c}d_{c}^{2}}$ (3)
where $d_{c}$ represents the number of samples selected per class, $n_{c}$ is the number of classes in the training dataset, and $\|\cdot\|_{2}$ denotes the Frobenius norm. We used $d_{c}$ = 16 for all the MedMNIST dataset experiments. The matrix $R_{C}$ encodes the pairwise similarities between the quantum representations of data samples, while $R_{\text{ref}}$ is a reference matrix capturing the ideal behavior of the circuit, where $R_{\text{ref}}(i,j)=1$ if two samples belong to the same class and 00 otherwise. A higher RepCap value suggests that the circuit has a greater capacity to distinguish between different classes in the dataset. Finally, a composite scoring metric ($F_{score}$) is computed based on CNR and RepCap to effectively select the most suitable quantum circuit defined as:
$F_{score}(C)=CNR(C)^{\alpha_{\text{CNR}}}\times RepCap(C)$ (4)
where, $\alpha_{\text{CNR}}$ is a tunable hyperparameter that controls the relative weight of noise resilience in the scoring process and we set $\alpha_{\text{CNR}}=0.5$ in our MedMNIST experiments.
@@ 2.3 Training quantum model on classical hardware
The training phase of our QML framework for medical image classification is executed on classical hardware as illustrated in Fig. 1. This approach circumvents the current limitations of quantum hardware in the NISQ era, such as limited qubits, high noise levels, and decoherence effects by performing the computationally expensive training process on classical GPUs even for large-scale datasets. Our QML workflow integrates classical optimization techniques with noise-resilient, hardware-efficient quantum circuits tailored for medical imaging tasks. The training pipeline includes iterative optimization using a gradient-based method and loss evaluation. The variational quantum circuit parameters $(\theta)$ are iteratively updated within a classical optimization loop to minimize a task-specific loss function. This iterative process continues until convergence is achieved or the maximum number of training epochs is reached, ensuring the parameters are optimized for the target classification task.
We conducted the training on noiseless simulators using the TorchQuantum library [41], which provides an ideal environment for parameter optimization without interference from hardware-induced noise. This approach significantly reduces computational overhead compared to direct hardware training and ensures consistent gradient computation. During each training iteration, the preprocessed input data is mapped to quantum states via encoding gates in the quantum circuit, which is then executed on the simulator. Expectation values obtained from circuit measurements are used to compute the loss function. Gradients of the loss function are calculated using backpropagation. Afterward, these gradients guide precise updates to the trainable parameters $(\theta)$. In our implementation, all the quantum circuits were trained for 200 epochs using the Adam optimizer with a learning rate of 0.01 and a batch size of 128, ensuring convergence to optimal parameter configurations. The MedMNIST contains datasets having a different number of classes. During training, we measure one qubit and two qubits for two class and four class datasets, respectively and a mean square loss function is used to calculate the loss. For the remaining multi-class datasets, we measure the number of qubits equal to the number of classes in the dataset, and the cross-entropy loss function is used to calculate the loss. To ensure compatibility with MedMNIST datasets, we trained circuits with varying numbers of trainable parameters, such as 60, 80, 100, and 120 for each dataset of MedMNIST to account for differences in dataset complexity. The best-performing models are saved along with their parameters, losses, and accuracies for later inference on real quantum hardware. We also performed inference of these trained quantum models on a simulator for all considered MedMNIST datasets. All quantum models are trained on an A100 Nvidia GPU with 43 GB RAM.